Assume that the salespeople of Excellent Corp. are paid 4% of sales as commission. The level of sales is $230,000 for this month, and the only fixed selling costs are $30,000. Excellent's budgeted selling costs for this month are:
Answer #1

Selling cost = Fixed cost + Variable Cost

selling Cost = 30,000 + ( 4% on 230,000)

= 30,000 + 9,200 = $39,200 (Answer)